Factor completely.
x^2 - 13x + 36
I get (x - 9) (x-4) by looking at it - but to break it down I am not sure how to get to this answer.
Thanks

Answer by mszlmb(115)

x^2 - 13x + 36

what two numbers add up to -13 and multiply to 36?
We know they are both negative for producing a positive product and a negative sum.
36=18*2=9*4 aha! sum of 13!
so..
x2-9x-4x+36 factor like terms
x(x-9)-4(x-9) put like terms together
(x-4)(x-9) ther ya go!
ax 4 questionz